The length of a rectangle is 32 units greater than its width. If the width is w, which expression gives the perimeter of the rectangle in terms of its width?
Answer by macston(5194)

Length=L=W+32: Width=W; Perimeter=P
P=2(L+W)
P=2(W+32+W)
P=2(2W+32)
P=4W+64